### v1

# Cascaded Self-Evaluation Augmented Training for Efficient Multimodal Large Language Models

Abstract
Efficient Multimodal Large Language Models (EMLLMs) have witnessed rapid development in recent years. Introducing Chain-of-Thought (CoT) during inference and leveraging self-evaluation step-by-step have further enhanced their performance. However, due to limited amount of the parameters, EMLLMs often perform poorly when directly using self-evaluation during inference. How to synthesize evaluation data? How much data to synthesize? What are the optimal training and inference strategies? What prompts to use? These are unexplored yet essential questions that need to be addressed to enhance the self-evaluation capabilities of EMLLMs.
To this end, we propose a method called Self-Evaluation Augmented Training (SEAT). We utilize more powerful EMLLMs to perform CoT reasoning, select beneficial data, and generate evaluations. The synthesized data is then used to train the EMLLMs. However, EMLLMs can’t handle long prompt inputs, and long CoT inference and self-evaluation outputs. Furthermore, synthesized data significantly diminishes its original CoT reasoning capabilities. To overcome these limitations, we further introduce Cascaded Self-Evaluation Augmented Training (Cas-SEAT). Cas-SEAT splits the long integrated prompts, which combine CoT reasoning and self-evaluation functionalities, into shorter cascaded prompts, each dedicated to one of these tasks. Additionally, we minimize the cost of Cas-SEAT to ensure its applicability in resource-constrained environments. During the synthesis of evaluation data, we use open-source EMLLMs with only 7B parameters and annotate a small amount of data using short prompts. Experiments show that, compared to existing methods, Cas-SEAT significantly enhances the self-evaluation capabilities of EMLLMs. It improves 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively. Additionally, as the first work to explore self-evaluation augmented training for EMLLMs, the Cas-SEAT Dataset we constructe to enhance the self-evaluation capabilities of EMLLMs also provides valuable resources and references for future research in this field.
1 Introduction
Figure 1: Several modes of self-evaluation: (a) A multimodal dataset containing images, questions, and answers. (b) The typical approach in existing methods, where the model performs self-evaluation directly during inference. (c) The primary model synthesizes self-evaluation data and uses it for its own training. (d) A more powerful model evaluates the primary model’s outputs, and the primary model learns from this feedback to develop self-evaluation capabilities. (e) A more powerful model evaluates a small subset of the primary model’s outputs, combines the evaluations with CoT data, and feeds them into the primary model. This approach also decomposes prompts and tasks to enable cascading self-evaluation capabilities.
In recent years, multimodal large language models (MLLMs) Liu et al. (2024); Wang et al. (2024) have advanced rapidly, achieving remarkable performance. However, broader training and deployment demands often come from resource-constrained settings, such as universities, hospitals, and communities. In these scenarios, the application of larger models, such as 72B or 40B-parameter MLLMs, faces significant limitations. This has spurred the emergence of efficient multimodal large language models (EMLLMs) with smaller parameter sizes, such as 7B, 2B, or even 1B. Today, EMLLMs are approaching the performance of early 40B-parameter MLLMs, thanks to advancements in model architecture, training methods, and techniques like Chain-of-Thought (CoT) Wei et al. (2022c). CoT, in particular, enables EMLLMs to explicitly model intermediate reasoning steps, producing more interpretable, logically coherent, and contextually relevant responses across a wide range of tasks involving text, images, and other modalities.
Despite these advancements, the further improvement of EMLLMs faces two major bottlenecks. First, the CoT data available for training is insufficient in both quantity and quality. Existing data, often sourced from web crawls or generated by models without strict quality control, suffers from inconsistencies, incomplete reasoning steps, logical gaps, and even incorrect reasoning. Such flawed samples not only fail to enhance model performance but may also degrade it. Second, while recent research suggests that self-evaluation can improve the accuracy and reliability of reasoning, these efforts primarily focus on refining prompts during the inference stage Xu et al. (2024); Lightman et al. (2024). They guide models to evaluate each reasoning step for accuracy but rarely generate evaluation data during training to enhance the self-evaluation capabilities of LLMs and MLLMs. Research on EMLLMs in this area remains non-existent. Due to their smaller parameter sizes, EMLLMs struggle to evaluate their own reasoning steps effectively or correct errors (as evidenced by the negligible performance improvement after evaluation in Figure 3’s black bar).
To address these challenges, we propose Self-Evaluation Augmented Training (SEAT). SEAT leverages more powerful EMLLMs to perform CoT reasoning and evaluate it. By refining these evaluations, we generate high-quality reasoning samples that help improve the overall reasoning quality and self-evaluation capabilities of the target model (referred to as the "main model" in this paper). However, the limited parameter scale of EMLLMs introduces two challenges to this process: (1) integrating reasoning and evaluation tasks into a single prompt leads to excessively long prompts during both input and output stages, overwhelming EMLLMs. For EMLLMs to understand that they need to perform CoT inference and self-evaluation, a long prompt is required. The output, consisting of step-by-step reasoning and the corresponding self-evaluation for each step, has to become even longer. (2) introducing extended evaluation data into EMLLMs can diminish their inherent CoT reasoning capabilities (as evidenced by SEAT’s poor inference and evaluation performance in Table  1, 2, 3, and 4).
To overcome these limitations, we further propose the Cascaded Self-Evaluation Augmented Training (Cas-SEAT) method. Cas-SEAT decouples the reasoning and self-evaluation processes into two separate tasks using two cascaded short prompts instead of a single lengthy integrated prompt. Regarding data handling, Cas-SEAT annotates only a small subset of data using the evaluation prompt in the evaluation phase and mixes this data with the reasoning prompt annotations from the reasoning phase. This approach ensures that EMLLMs first enhance their CoT reasoning capabilities before learning self-evaluation, as we believe that robust CoT reasoning is the foundation for effective self-evaluation. We also minimize the cost of Cas-SEAT to ensure its applicability in resource-constrained environments. During evaluation data synthesis, we use only open-source EMLLMs with 7B parameters and annotate a small dataset with short prompts. This method primarily generates high-quality reasoning samples, adding minimal self-evaluation data to optimize both data utility and training efficiency. Experimental results show that Cas-SEAT significantly improves self-evaluation capabilities compared to existing methods, with performance gains of 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively. Equally important, as the first work to explore self-evaluation augmented training for EMLLMs, both Cas-SEAT and the Cas-SEAT Dataset we constructed provide a valuable foundation and reference for future research in this domain.
In summary, our contributions are as follows:
  • We propose SEAT, exploring how EMLLMs improve their CoT reasoning and self-evaluation capabilities by learning from CoT-generated self-evaluation data. To the best of our knowledge, this is the first study of its kind on EMLLMs and is rare even among LLMs and MLLMs.
  • We designed Cas-SEAT to effectively address the challenges EMLLMs face with handling long prompt inputs and lengthy CoT and self-evaluation outputs, successfully enhancing self-evaluation capabilities while maintaining CoT reasoning abilities.
  • We construct the Cas-SEAT Dataset, the first dataset tailored for self-evaluation augmented training of EMLLMs. It is low-cost yet effective, facilitating future research.
  • We conducted extensive experiments across multiple datasets, exploring different model architectures (LLaVA-1.5, Qwen2-VL) and parameter scales (7B, 2B). We performed a detailed analysis of overall performance, variations in performance across different aspects, and further investigated the CoT reasoning ability and self-evaluation capability. The experimental results demonstrate that utilizing Cas-SEAT significantly enhances performance.
2 Related Work
CoT Reasoning.
CoT reasoning Huang et al. (2022); Hoffmann et al. (2022); Chowdhery et al. (2022); Wei et al. (2022a); Zhang et al. (2022); Kojima et al. (2022) enhances the performance of LLMs by forcing them to perform step-by-step reasoning. The initial idea was proposed by Wei et al. (2022b), demonstrating that simply modifying the prompt could significantly improve model performance on complex tasks. Following Wei et al. (2022b), numerous studies have shown that LLMs can perform CoT reasoning through zero-shot CoT prompting (Kojima et al. (2022)) and few-shot CoT prompting (Wei et al. (2022b); Zhang et al. (2022)). Improvements have been made in areas such as self-consistency (Wang et al., 2022b), least-to-most prompting (Zhou et al., 2022), dynamic least-to-most prompting (Drozdov et al., 2022), bootstrapping training (Zelikman et al., 2022), and self-training (Huang et al., 2022). Wang et al. (2022b, a) refined the original CoT prompting by marginalizing over diverse reasoning paths, while Zelikman et al. (2022); Huang et al. (2022) improved CoT prompting through bootstrapping on self-generated CoT prompts. Li et al. (2022a) introduced a voting classifier to filter sampled CoT prompts before the final prediction. Additionally, prompt enhancement and selection (Shum et al., 2023), meta-heuristic approaches (Pan et al., 2023), and meta-graph prompting (Pan et al., 2024) have further advanced standard CoT prompting. Existing CoT studies pay little attention to the correctness of synthesized CoT data, especially lacking focus on the CoT capabilities of EMLLMs. CoT reasoning is more challenging for EMLLMs. Our Cas-SEAT improves the quality of synthesized training data through self-evaluation and enhances the CoT reasoning capabilities of EMLLMs.
LLM Self-Evaluation.
Recent research on LLM calibration has shown that current LLMs can achieve well-calibrated predictions in specific tasks (Rae et al., 2021; Kadavath et al., 2022; Guo et al., 2017; Kadavath et al., 2022; Jiang et al., 2021; Kuhn et al., 2023; Fu et al., 2023). For instance, Liu et al. Liu et al. (2023a) leveraged CoT prompting to use GPT-4 for text evaluation. Kocmi and Federmann Kocmi and Federmann (2023); Xu et al. Xu et al. (2023) designed methods focused on enabling LLMs to produce more fine-grained correction annotations. Additionally, Koo et al. Koo et al. (2023), Zheng et al. Zheng et al. (2023), Chang et al. Chang et al. (2023), Deutsch et al. Deutsch et al. (2022), and Liu et al. Liu et al. (2023b) explored issues of self-reinforcing bias and fairness in LLMs. It is worth noting that scaling up model size plays a crucial role in improving calibration (Rae et al., 2021; Wei et al., 2022a). As LLMs exhibit strong calibration capabilities, more research has focused on prompting LLMs for self-evaluation (Golovneva et al., 2022; Zhang et al., 2023; Shinn et al., 2023; Madaan et al., 2024; Paul et al., 2023). Although dedicated evaluators have proven effective in assessing the correctness of reasoning data (Li et al., 2022b; Xu et al., 2024), self-evaluation is more efficient compared to using external models for evaluation, as it eliminates the need for additional annotations and evaluators (Li et al., 2022a). Unlike existing methods that optimize the reasoning process through self-evaluation, our approach uses synthetic self-evaluation data to enhance model training. Moreover, we focus on EMLLMs, which introduce additional challenges to the training and reasoning processes of self-evaluation.
3 Methodology
Figure 2: Overview of the method. It illustrates the prompts designed for Augmented Self-Evaluation and Augmented Cascading Self-Evaluation, along with the corresponding training data generated.
3.1 Problem Formulation and Notations
Data and Model.
The primary EMLLM model is denoted as $\mathcal{M}$, with parameters $\Theta_{0}$. To achieve better performance on the target task, $\mathcal{M}$ needs to be fine-tuned on the dataset $\mathcal{D}$. We use $\{\mathcal{X},\mathcal{Y}\}$ to represent a sample, $\mathcal{X}$ include images $\mathcal{I}$ and queries $\mathcal{Q}$, and $\mathcal{Y}$ represents answers. After fine-tuning, $\Theta_{0}$ will be optimized to $\Theta$. The output of $\mathcal{M}$ are denoted as $\hat{\mathcal{Y}}$. $\mathcal{D}_{\rm{se}}$ represents the self-evaluation data generated by $\mathcal{M}$. A more powerful EMLLM model is denoted as $\mathcal{G}$, which generates CoT reasoning data based on $\mathcal{D}$, denoted as $\mathcal{S}_{\rm{cot}}$. After evaluating $\mathcal{S}_{\rm{cot}}$, $\mathcal{G}$ produces self-evaluation training data, denoted as $\mathcal{S}_{\rm{se}}$. A small subset of $\mathcal{S}_{\rm{cot}}$ is selected, denoted as $\mathcal{S}_{\rm{cot}}^{\rm{sub}}$, and after being evaluated by $\mathcal{G}$, generates cascaded self-evaluation training data, denoted as $\mathcal{S}_{\rm{cse}}$.
Prompt.
The prompts used for $\mathcal{M}$ inference and self-evaluation are denoted as $P$ and $P_{se}$, respectively. The prompts used by $\mathcal{G}$ to generate cascading CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{se}$, respectively. The prompts used by $\mathcal{G}$ to generate CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{cse}$, respectively. Identical symbols indicate identical prompts.
Formula.
We formalize the existing studies to facilitate a formula-based comparison of their differences. $l$ represents the loss function.
Inference:
$\hat{\mathcal{Y}}=\mathcal{M}(\{P\!\!\quad\text{or}\!\!\quad P_{\rm{cot}}\!\!\quad\text{or}\!\!\quad P_{\rm{se}}\},\mathcal{X};\Theta_{0}),\mathcal{X}\in\mathcal{D}$
Finetune:
$\operatorname*{arg\,min}\limits_{\Theta_{0}}\mathcal{L}=\sum\limits_{\mathcal{X},\mathcal{Y}\in\mathcal{D}}l(\mathcal{Y},\mathcal{M}(P\!\!\quad\text{or}\!\!\quad P_{\rm{cot}},\mathcal{X};\Theta_{0}))$
Augmented CoT:
$&\mathcal{S}_{\rm{cot}}=\mathcal{G}(P_{\rm{cot}},\mathcal{X};\Theta)\\ &\operatorname*{arg\,min}_{\Theta}\mathcal{L}=\sum_{\mathcal{X},\mathcal{Y}\in\mathcal{S}_{\rm{cot}}}l(\mathcal{Y},\mathcal{M}(P\!\!\quad\text{or}\!\!\quad P_{\rm{cot}},\mathcal{X};\Theta))\end{aligned}\right.$
SEAT:
[…]
Cas-SEAT:
[…]
3.2 SEAT
Figure 3: Comparison of evaluation in terms of model performance improvement
Since self-evaluation has been proven effective in LLMs and MLLMs, we tested the self-evaluation performance of EMLLM, as shown in Figure 3. Specifically, we evaluated inference and post-inference evaluation using the LLaVA-based Base(pretrained model), Finetune (finetune on Math360k), and Augmented CoT (Qwen2-VL (7B) annotates CoT data on Math360k, which is then learned by LLaVA-1.5.) on MathVista, Math-V, and We-Math. We observed that, in most cases, evaluation led to some improvement, though the improvement was not significant. However, in certain cases, evaluation resulted in a performance decline. We believe this variability in performance is primarily due to the lack of evaluation-specific data during training. Therefore, we aim to enhance the evaluation capabilities of EMLLM by synthesizing more self-evaluation data for its training.
The prompt $P_{\rm{se}}$ and the training data $\mathcal{S}_{\rm{se}}$ for Augmented Self-Evaluation are shown in Figure 2. The prompt divides Augmented Self-Evaluation into two parts: CoT reasoning and self-evaluation. The self-evaluation also includes data selection, where the MLLM autonomously determines which data to use for learning and which to discard. The learning process can be formulated as Equation 4.
3.3 Cas-SEAT
Due to the parameter limitations of EMLLM, its performance significantly degrades in scenarios involving long-token inputs and outputs. Consequently, Augmented Self-Evaluation often leads to a decline in performance. To address this issue, we decouple the inference and evaluation processes, as illustrated in Figure 2. First, we use a more powerful EMLLM $\mathcal{G}$ to obtain $\mathcal{S}_{\rm{cot}}$ based on $P_{\rm{cot}}$. Next, we continue to use $\mathcal{G}$ to evaluate the incorrect parts of $\mathcal{S}_{\rm{cot}}$, denoted as $Error(\mathcal{S}_{\rm{cot}})$. This process produces $\mathcal{S}_{\rm{cse}}=\mathcal{G}(P_{\rm{cse}},Error(\mathcal{S}_{\rm{cot}}))$. Subsequently, we retain the data corrected through evaluation, denoted as $Correct(\mathcal{S}_{\rm{cse}})$, which satisfies the relationship $Correct(\mathcal{S}_{\rm{cse}})\subset\mathcal{S}_{\rm{cse}}$. For simplicity, $Correct(\mathcal{S}_{\rm{cse}})$ is abbreviated as $\mathcal{S}_{\rm{cse}}$. So the whole equation to get $\mathcal{S}_{\rm{cse}}$ is:
[…]
In the above equation, $A:=B$ means assigning the value of $B$ to $A$. Then, $\mathcal{M}$ is fine-tuned on $\mathcal{S}_{\rm{cse}}$ like Equation 5.
4 Experiments
We conducted experiments to evaluate the effectiveness and generalizability of Cas-SEAT.
4.1 Experimental Setup
4.1.1 Datasets
We train on MathV360K dataset and evaluate on Math-Vista, Math-V, We-Math. MathV360K is a widely used dataset for MLLM training. Math-Vista), Math-V, We-Math are three widely used public benchmarks for MLLM evaluation.
MathVista
Method Average Multi-choice Free-form Text Integer General VQA Math-targeted VQA
Base (inference) 0.2850 0.4407 0.1022 0.4407 0.1124 0.3391 0.2389
[…]
Finetune (inference) 0.3160 0.4481 0.1609 0.4481 0.1770 0.3196 0.3130
Finetune (evaluation) 0.3490 0.4926 0.1804 0.4926 0.1986 0.3543 0.3444
CoT (inference) 0.3380 0.4815 0.1696 0.4815 0.1866 0.3326 0.3426
[…]
SEAT (inference) 0.2760 0.4278 0.0978 0.4278 0.1077 0.2957 0.2593
SEAT (evaluation) 0.2850 0.4389 0.1043 0.4389 0.1148 0.3196 0.2556
Cas-SEAT (inference) 0.3390 0.4889 0.1630 0.4889 0.1794 0.3652 0.3167
Cas-SEAT (evaluation) 0.4500 0.6222 0.2478 0.6222 0.2727 0.4848 0.4204
Improve 19.68% 15.07% 31.04% 15.07% 31.04% 11.50% 17.01%
Table 1: Comparison of Cas-SEAT and the Baseline on MathVista Dataset
Math-V
Method All Level1 Level2 Level3 Level4 Level5 ALG ARI CG COM
Base (inference) 0.0526 0.0800 0.0690 0.0364 0.0444 0.0299 0.0000 0.0000 0.2941 0.0526
[…]
SEAT (inference) 0.0592 0.1132 0.0488 0.0357 0.0000 0.0882 $0$ul0.0526 0.1579 0.0000
SEAT (evaluation) 0.0888 0.1509 0.0732 0.0714 0.0000 0.1324 $0$ul0.0526 0.1579 0.0526
Cas-SEAT (inference) 0.1711 0.1321 0.1341 0.2321 0.1778 0.1912 0.1579 0.1579 0.2632 0.1579
Cas-SEAT (evaluation) 0.2763 0.2642 0.2439 0.2500 0.3111 0.3235 0.3158 0.1579 0.4211 0.2105
Improve 55.57% 27.33% 25.01% 179.96% 40.01% 57.12% 199.91% 200.19% 14.31% 99.91%
Table 2: Comparison of Cas-SEAT and the Baseline on Math-V Dataset
We-Math
Method Avg(Strict) IG(Strict) CM(Strict) Avg(Loose) IG(Loose) CM(Loose)
Base (inference) 0.0143 0.0171 0.0057 0.0657 0.0171 0.0571
[…]
Finetune (inference) 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
Finetune (evaluation) 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
CoT (inference) 0.0438 0.0305 0.0286 0.1752 0.0305 0.1600
[…]
SEAT (inference) 0.0105 0.0095 0.0057 0.0429 0.0095 0.0381
SEAT (evaluation) 0.0114 0.0114 0.0057 0.0495 0.0114 0.0438
Cas-SEAT (inference) 0.0533 0.0343 0.0362 0.2114 0.0343 0.1943
Cas-SEAT (evaluation) 0.1076 0.0438 0.0857 0.3495 0.0438 0.3276
Improve 46.79% 9.50% 55.25% 23.54% 9.50% 24.61%
Table 3: Comparison of Cas-SEAT and the baselines on We-Math dataset
4.1.2 Baselines
To verify the applicability, the following EMLLMs are implemented and compared with the counterparts combined with the proposed method. We primarily analyzed the effectiveness of our method based on the EMLLMs LLaVA-v1.5(7B) Liu et al. (2024), Qwen2-VL(2B) Wang et al. (2024). Since current EMLLMs research often involves training on various datasets, including those frequently used as test datasets, such as the ones we selected, and many MLLM studies do not disclose their training datasets, we opted not to include EMLLMs published after the release of these datasets to ensure a fair comparison.
4.1.3 Implementation Details
“Base”: The pretrained MLLM Liu et al. (2024); Wang et al. (2024). “Finetune”: The pretrained MLLM is lora fine-tuned on Math360k Hu et al. (2022). “CoT”: Qwen2-VL (7B) performs CoT reasoning on Math360k. The MLLM then learns from this data Wang et al. (2023); Wadhwa et al. (2024). “SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k. Qwen2-VL (7B) evaluates the reasoning outputs of the MLLM, and the MLLM learns from this evaluated data Xiong et al. (2024). “Cas-SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k. Qwen2-VL (7B) evaluates the MLLM’s incorrect reasoning outputs. These evaluated data are combined with CoT data, and the MLLM learns from this combined dataset.
We-Math
Method UCU AL CPF UPF CSF USF BTF One-step Two-step Three-step
[…]
Finetune (inference) 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.1950 0.3004 0.2750 0.3394
Finetune (evaluation) 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.2172 0.3021 0.2750 0.3394
[…]
SEAT (inference) 0.0942 0.0263 0.1099 0.1476 0.1083 0.1624 0.1176 0.1210 0.1361 0.1091
SEAT (evaluation) 0.0942 0.0263 0.1179 0.1476 0.1130 0.1624 0.1176 0.1243 0.1472 0.1091
Cas-SEAT (inference) 0.2679 0.3439 0.3240 0.3216 0.3230 0.2408 0.3240 0.3103 0.2861 0.2364
Cas-SEAT (evaluation) 0.4256 0.1982 0.4294 0.3546 0.3303 0.3431 0.3462 0.3909 0.3528 0.4788
Improve 8.32% -14.42% 24.36% 16.76% 11.44% 4.60% 34.71% 20.54% 13.40% 27.41%
Table 4: Detailed comparison of Cas-SEAT and the baselines on We-Math dataset
After obtaining outputs on the test dataset, we first use a more powerful MLLM to extract answers for the MathVista dataset, followed by using regular expressions to calculate accuracy. For Math-V and We-Math, we directly use the regular expressions provided in the datasets to calculate accuracy.
4.2 Experimental Results
4.2.1 Overall Assessment of Cas-SEAT
As shown in Tables 1, 2, 3, and 4, we analyzed the experimental results on the MathVista, Math-V, and We-Math datasets. We evaluated the accuracy of two types of outputs for each method: direct inference (inference) and self-evaluation (evaluation). In addition to overall performance, we conducted tests on multiple subsets of the three datasets, including generalization capability, data types, problem difficulty, and task types. Below is the correspondence between these categories and the columns in the tables. (Due to dataset limitations, some categories could not be tested for certain datasets.): Overall Performance: MathVista (Table 1: Average), Math-V (Table 2: All), We-Math (Table 3: Score (Strict), Score (Loose)). Cas-SEAT shows very significant improvements over the baseline on all datasets, achieving around a 50% improvement on both Math-V and We-Math. Generalization Capability: We-Math (Table 3: Inadequate Generalization (Strict), Complete Mastery (Strict), Inadequate Generalization (Loose), Complete Mastery (Loose)). The improvement range of 9.5% to 55.25% demonstrates Cas-SEAT’s strong generalization capabilities, excelling at problems involving more integrated knowledge points, as evidenced by the significant improvement in Complete Mastery (CM). Data Types: MathVista (Table 1: Text, Integer). Cas-SEAT is better at answering integer-type questions. Problem Difficulty: MathVista (Table 1: Multi-choice, Free-form), Math-V (Table 2: Level1, Level2, Level3, Level4, Level5), We-Math (Table 3: One-step, Two-step, Three-step). The Math-V subsets are divided into five difficulty levels, with higher levels indicating more complex problems. The We-Math subsets are classified according to the number of reasoning steps, with more steps indicating higher difficulty. Cas-SEAT shows a more pronounced advantage on more challenging problems. Task Types: MathVista: General VQA, Math-targeted VQA; Math-V: Algebra (ALG), Arithmetic (ARI), Combinatorial Geometry (CG), Combinatorics (COM); We-Math: Understanding and Conversion of Units (UCU), Angles and Length (AL), Calculation of Plane Figures (CPF), Understanding of Plane Figures (UPF), Calculation of Solid Figures (CSF), Understanding of Solid Figures (USF), Basic Transformations of Figures (BTF). Cas-SEAT exhibits significant improvements over the best baseline method on almost all problem types, especially for numerical computation tasks (Math-V: ALG, ARI). For VQA tasks, Cas-SEAT performs better on the more difficult Math-targeted VQA.
In summary, we found that: 1. After fine-tuning, the inference ability of EMLLMs improves, but their self-evaluation ability declines. This is because fine-tuning enhances their mathematical reasoning ability but leads to forgetting general knowledge. Using synthetic CoT data based on mathematical datasets for enhanced training shows a similar pattern. 2. SEAT, which integrates CoT inference and self-evaluation training, suffers a significant drop in CoT inference ability, without a substantial improvement in self-evaluation. Even if EMLLMs are first trained with synthetic CoT data, using only self-evaluation training in the second round significantly reduces their CoT reasoning ability. This demonstrates the detrimental effect of long prompts and outputs on EMLLMs, and the difficulty of maintaining both CoT inference and self-evaluation capabilities simultaneously. 3. Cas-SEAT, by separating the CoT inference and self-evaluation tasks and shortening the prompt, preserves robust inference ability. More importantly, it achieves far stronger self-evaluation performance than the baselines. Our approach Cas-SEAT achieves remarkably significant improvements over the best baseline methods, both overall and across each subset. The detailed analysis shows that Cas-SEAT is especially adept at more difficult numerical computation tasks.
4.2.2 Evaluation on Different EMLLMs
Figure 4: Comparison of Cas-SEAT and the Baseline based on Qwen2-VL(2B).
Figure 5: A bar chart analysis sample in MathVista. Green background indicates the raw data, red text represents incorrect reasoning processes (sometimes with no reasoning process), pink background and yellow background denote results from direct reasoning and self-evaluation, respectively. Blue text and blue background indicate the corrected reasoning process and corrected results, respectively.
As shown Figure 4, we used Qwen2-VL (2B) on Math-V to compare the performance of Cas-SEAT with that of the baseline. We take the maximum value of all methods as the circumference and 0 as the center of the circle. The results demonstrated a significant advantage for Cas-SEAT over the baselines. This indicates that our approach is also applicable to smaller EMLLMs with different architectures.
4.2.3 Case Study
We present two representative examples in Figure 5 and Figure 6 (in appendix). From Figure 5 and Figure 6, the following observations can be made: 1. Reasoning Process and Its Accuracy: Direct reasoning using various methods failed to produce the correct answer. The Base, Finetune, and SEAT methods generally lack a complete reasoning process and jump directly to the final answer. Incorporating CoT data into EMLLM training (e.g., CoT, Cas-SEAT) enables step-by-step reasoning, but it does not always guarantee accurate results. Moreover, although self-evaluation data is constructed based on CoT data, the lengthy prompts and evaluation data diminish the model’s reasoning capability. 2. Evaluation Process and Its Accuracy: Generating self-evaluation data from CoT data and applying it to training seems to enhance EMLLM’s self-evaluation capability by enabling targeted training to help the model reflect on whether its outputs are reasonable. However, in many cases, such reflections are superficial, and a large number of self-evaluation instances are ineffective. In contrast, Cas-SEAT effectively preserves the model’s reasoning ability while significantly enhancing its self-evaluation capability. By mixing a small amount of self-evaluation data into the CoT data, EMLLM maintains its CoT reasoning ability while upgrading its self-evaluation skills. These conclusions hold consistently across several datasets.
5 Conclusion
We proposed SEAT and its enhanced variant, Cas-SEAT, to improve the self-evaluation capabilities of Efficient Multimodal Large Language Models (EMLLMs). By synthesizing evaluation data with stronger models and using cascaded task decomposition, Cas-SEAT enhances performance while balancing CoT reasoning and self-evaluation, even under resource constraints. Experiments show significant gains on benchmark datasets, and the Cas-SEAT Dataset provides a valuable resource for future research, laying a strong foundation for advancing self-evaluation in EMLLMs.
Limitation
Potential limitations may include a small amount of resource consumption. The method is designed specifically for EMLLMs, and it may not currently be the optimal solution for larger MLLMs.
Appendix A Appendix
A.1 Supplementary Experiments
A.1.1 Case Study
Figure 6: A table analysis sample in MathVista. The meanings of the colors are the same as those in Figure 5.
Figure 6 is a supplementary case, the observed findings can be referred to the section 4.2.3.
A.1.2 Datasets
The statistics of the training dataset used in the experiments is shown in Table 5.
Table 5: Statistics of Datasets.
Dataset MathV360k CoT Cas-SEAT
#Samples 339k 160k 167k
A.1.3 Hyperparameters and Training Schedules
We summarize the hyperparameters and training schedules of the LLaVA-1.5 and Qwen2-VL used in the experiments. Table 6 shows the settings of the LLaVA-1.5 training. Table 7 shows the settings of the Qwen2-VL training.
Table 6: Hyperparameters and training schedules of LLaVA.
MLLM Hyperparameter Setting
LLaVA-1.5- 7-hf GPU Tesla A100 (40GB)
Batch Size 20
LoRA Rank 128
LoRA Alpha 64
LoRA Dropout 0
Optimizer AdamW
Warmup Steps 50
Learning tate 2e-5
Epochs 2
Table 7: Hyperparameters and training schedules of Qwen2-VL.
MLLM Hyperparameter Setting
Qwen2-VL- 2B-Instruct GPU Tesla A100 (40GB)
Batch Size 20
LoRA Rank 128
LoRA Alpha 64
LoRA Dropout 0
Optimizer AdamW
Warmup Steps 50
Learning tate 2e-5
Epochs 2

### v2

@@
-Cascaded Self-Evaluation Augmented Training for Efficient Multimodal Large Language Models
-Abstract
-Efficient Multimodal Large Language Models (EMLLMs) have witnessed rapid development in recent years.
-Introducing Chain-of-Thought (CoT) during inference and leveraging self-evaluation step-by-step have further enhanced their performance.
-However, due to limited amount of the parameters, EMLLMs often perform poorly when directly using self-evaluation during inference.
-How to synthesize evaluation data?
-How much data to synthesize?
-What are the optimal training and inference strategies?
-What prompts to use?
-These are unexplored yet essential questions that need to be addressed to enhance the self-evaluation capabilities of EMLLMs.
-To this end, we propose a method called Self-Evaluation Augmented Training (SEAT).
-We utilize more powerful EMLLMs to perform CoT reasoning, select beneficial data, and generate evaluations.
-The synthesized data is then used to train the EMLLMs.
-However, EMLLMs can’t handle long prompt inputs, and long CoT inference and self-evaluation outputs.
-Furthermore, synthesized data significantly diminishes its original CoT reasoning capabilities.
-To overcome these limitations, we further introduce Cascaded Self-Evaluation Augmented Training (Cas-SEAT).
-Cas-SEAT splits the long integrated prompts, which combine CoT reasoning and self-evaluation functionalities, into shorter cascaded prompts, each dedicated to one of these tasks.
-Additionally, we minimize the cost of Cas-SEAT to ensure its applicability in resource-constrained environments.
-During the synthesis of evaluation data, we use open-source EMLLMs with only 7B parameters and annotate a small amount of data using short prompts.
-Experiments show that, compared to existing methods, Cas-SEAT significantly enhances the self-evaluation capabilities of EMLLMs.
-It improves 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively.
-Additionally, as the first work to explore self-evaluation augmented training for EMLLMs, the Cas-SEAT Dataset we constructe to enhance the self-evaluation capabilities of EMLLMs also provides valuable resources and references for future research in this field.
-1 Introduction
-Figure 1: Several modes of self-evaluation: (a) A multimodal dataset containing images, questions, and answers.
-(b) The typical approach in existing methods, where the model performs self-evaluation directly during inference.
-(c) The primary model synthesizes self-evaluation data and uses it for its own training.
-(d) A more powerful model evaluates the primary model’s outputs, and the primary model learns from this feedback to develop self-evaluation capabilities.
-(e) A more powerful model evaluates a small subset of the primary model’s outputs, combines the evaluations with CoT data, and feeds them into the primary model.
-This approach also decomposes prompts and tasks to enable cascading self-evaluation capabilities.
-In recent years, multimodal large language models (MLLMs) Liu et al.
-(2024); Wang et al.
-(2024) have advanced rapidly, achieving remarkable performance.
-However, broader training and deployment demands often come from resource-constrained settings, such as universities, hospitals, and communities.
-In these scenarios, the application of larger models, such as 72B or 40B-parameter MLLMs, faces significant limitations.
-This has spurred the emergence of efficient multimodal large language models (EMLLMs) with smaller parameter sizes, such as 7B, 2B, or even 1B.
-Today, EMLLMs are approaching the performance of early 40B-parameter MLLMs, thanks to advancements in model architecture, training methods, and techniques like Chain-of-Thought (CoT) Wei et al.
-(2022c).
-CoT, in particular, enables EMLLMs to explicitly model intermediate reasoning steps, producing more interpretable, logically coherent, and contextually relevant responses across a wide range of tasks involving text, images, and other modalities.
-Despite these advancements, the further improvement of EMLLMs faces two major bottlenecks.
-First, the CoT data available for training is insufficient in both quantity and quality.
-Existing data, often sourced from web crawls or generated by models without strict quality control, suffers from inconsistencies, incomplete reasoning steps, logical gaps, and even incorrect reasoning.
-Such flawed samples not only fail to enhance model performance but may also degrade it.
-Second, while recent research suggests that self-evaluation can improve the accuracy and reliability of reasoning, these efforts primarily focus on refining prompts during the inference stage Xu et al.
-(2024); Lightman et al.
-(2024).
-They guide models to evaluate each reasoning step for accuracy but rarely generate evaluation data during training to enhance the self-evaluation capabilities of LLMs and MLLMs.
-Research on EMLLMs in this area remains non-existent.
-Due to their smaller parameter sizes, EMLLMs struggle to evaluate their own reasoning steps effectively or correct errors (as evidenced by the negligible performance improvement after evaluation in Figure 3’s black bar).
-To address these challenges, we propose Self-Evaluation Augmented Training (SEAT).
-SEAT leverages more powerful EMLLMs to perform CoT reasoning and evaluate it.
-By refining these evaluations, we generate high-quality reasoning samples that help improve the overall reasoning quality and self-evaluation capabilities of the target model (referred to as the "main model" in this paper).
-However, the limited parameter scale of EMLLMs introduces two challenges to this process: (1) integrating reasoning and evaluation tasks into a single prompt leads to excessively long prompts during both input and output stages, overwhelming EMLLMs.
-For EMLLMs to understand that they need to perform CoT inference and self-evaluation, a long prompt is required.
-The output, consisting of step-by-step reasoning and the corresponding self-evaluation for each step, has to become even longer.
-(2) introducing extended evaluation data into EMLLMs can diminish their inherent CoT reasoning capabilities (as evidenced by SEAT’s poor inference and evaluation performance in Table  1, 2, 3, and 4).
-To overcome these limitations, we further propose the Cascaded Self-Evaluation Augmented Training (Cas-SEAT) method.
-Cas-SEAT decouples the reasoning and self-evaluation processes into two separate tasks using two cascaded short prompts instead of a single lengthy integrated prompt.
-Regarding data handling, Cas-SEAT annotates only a small subset of data using the evaluation prompt in the evaluation phase and mixes this data with the reasoning prompt annotations from the reasoning phase.
-This approach ensures that EMLLMs first enhance their CoT reasoning capabilities before learning self-evaluation, as we believe that robust CoT reasoning is the foundation for effective self-evaluation.
-We also minimize the cost of Cas-SEAT to ensure its applicability in resource-constrained environments.
-During evaluation data synthesis, we use only open-source EMLLMs with 7B parameters and annotate a small dataset with short prompts.
-This method primarily generates high-quality reasoning samples, adding minimal self-evaluation data to optimize both data utility and training efficiency.
-Experimental results show that Cas-SEAT significantly improves self-evaluation capabilities compared to existing methods, with performance gains of 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively.
-Equally important, as the first work to explore self-evaluation augmented training for EMLLMs, both Cas-SEAT and the Cas-SEAT Dataset we constructed provide a valuable foundation and reference for future research in this domain.
+Cascaded Self-Evaluation Augmented Training for Lightweight Multimodal LLMs
+Abstract.
+Efficient Multimodal Large Language Models (EMLLMs) can improve performance through Chain-of-Thought (CoT) reasoning, but they have poor self-evaluation capabilities during the CoT reasoning process.
+This is due to their tendency to simplify the reasoning process and the degradation of self-evaluation ability during downstream task fine-tuning.
+To address this, we intuitively propose Self-Evaluation Augmented Training (SEAT), which uses more powerful EMLLMs to evaluate CoT reasoning data.
+The evaluation data is then used to train EMLLMs.
+However, due to the difficulties EMLLMs face with processing long token input-output sequences, and the degradation of self-evaluation ability as a basis for CoT reasoning, the SEAT method is not fully adapted.
+Therefore, we further propose Cascaded Self-Evaluation Augmented Training (Cas-SEAT), which converts long prompts into cascaded short prompts, each focusing on a specific task.
+Additionally, we mix CoT reasoning and self-evaluation data to preserve its CoT reasoning ability while enhancing the self-evaluation capability of EMLLMs.
+We also conduct Double-level Data Filtering (DDF), which includes source data filtering and labeled data filtering, using both non-EMLLM and EMLLM for filtering.
+Cas-SEAT and DDF work together to improve the performance of EMLLMs.
+Experiments show that Cas-SEAT achieves an average improvement of 22.16% across multiple datasets, and DDF significantly reduces the resource consumption of training  [^1] The code and a portion of the Cas-SEAT dataset are available at https://github.com/HelloZicky/Cas-SEAT.
+Multimodal LLM, Self-evaluation, Chain-of-thought
+1. Introduction
+Figure 1.
+(a) A sample from the dataset used for training and inference of multimodal large language models, containing images, questions, and answers.
+(b) Overview of Chain-of-Thought (CoT) reasoning, self-evaluation reasoning, and their corresponding enhancement methods.
+(c) The proposed computational method, Cas-SEAT.
+(d) The proposed dataset construction method, DDF, which provides the Cas-SEAT Dataset for Cas-SEAT.
+(e) Comparison of CoT reasoning ability, self-evaluation ability, and overall performance.
+Symbols “–”, “↑”, “↓”, “↑↑”, and “↓↓” indicate comparable, improved, degraded, significantly improved, and significantly degraded performance, respectively.
+In recent years, multimodal large language models (MLLMs) (Liu et al., 2024a; Wang et al., 2024a) have developed rapidly.
+However, the growing demand for training and deployment in resource-constrained environments such as universities, hospitals, and communities has spurred interest in developing efficient multimodal large language models (EMLLMs) with smaller parameter sizes, such as 7B, 2B, or even 1B.
+Thanks to advances in model architecture, training methods, and techniques like chain-of-thought (CoT) reasoning (Wei et al., 2022c), EMLLMs can now generate step-by-step reasoning processes that are more logical and coherent.
+With the widespread use of CoT reasoning, the accuracy of each step in the reasoning process has received increasing attention (Madaan et al., 2024; Zheng et al., 2023; Liu et al., 2023b), which we refer to as self-evaluation.
+However, directly applying these techniques to EMLLMs has not been effective (as shown by the black bars in Figure 3, where the improvement is minimal).
+This is primarily due to two reasons: (1) EMLLMs tend to “cut corners” when answering questions, making the reasoning process very short.
+(2) Due to the limited parameter size of EMLLMs, their self-evaluation ability is weak unless enhanced training is performed, rather than just pretraining.
+More critically, before deploying EMLLMs in applications, traditional supervised fine-tuning (SFT) is often applied to downstream tasks, which further weakens their already fragile self-evaluation ability (as shown in Table 2, where fine-tuned EMLLMs perform worse in self-evaluation (Line 14) compared to pretrained models (Line 13)).
+To address these challenges, we intuitively propose a Self-Evaluation Augmented Training (SEAT) method to enhance the self-evaluation capability of EMLLMs.
+SEAT leverages more powerful EMLLMs to perform CoT reasoning and evaluate them.
+By optimizing these evaluations, we generate high-quality reasoning samples that help improve the overall reasoning quality and self-evaluation ability of the target model (referred to as the “main model” in this paper).
+However, the vanilla SEAT method also faces two challenges: (1) Merging reasoning and evaluation tasks into a single prompt leads to excessively long inputs and outputs.
+From the input perspective, a long prompt is required for EMLLMs to handle both step-by-step reasoning and self-evaluation.
+From the output perspective, EMLLMs need to perform step-by-step reasoning followed by step-by-step self-evaluation.
+However, EMLLMs struggle with handling long token inputs and outputs.
+(2) The CoT reasoning ability of EMLLMs significantly drops after adding self-evaluation data for enhanced training, and CoT reasoning is the foundation of self-evaluation (as shown in the SEAT reasoning and evaluation performance in Tables 1, 2, 3).
+To overcome these limitations, we further improved SEAT by proposing the Cascaded Self-Evaluation Augmented Training (Cas-SEAT) method.
+Cas-SEAT decouples the reasoning and self-evaluation processes into independent tasks, using a cascade of brief prompts, with each prompt focusing on a specific task.
+In order for Cas-SEAT to work, we primarily use labeled CoT data, mixed with a small amount of self-evaluation data.
+The mixed data is used for training the main model to ensure that the CoT reasoning ability is preserved while enhancing the self-evaluation capability of EMLLMs.
+Due to the weaker generalization ability of EMLLMs, many data beneficial for training large MLLMs are not suitable for EMLLMs, so we use Double-level Data Filtering (DDF) for data selection.
+DDF includes source data filtering and labeled data filtering, and the filtering methods involve both non-EMLLM and EMLLM filtering.
+DDF primarily considers selecting data beneficial for EMLLMs training from six aspects: image quality, text quality, text length, text format, problem domain, and problem difficulty.
+This ensures the model’s responses are standardized while avoiding data that exceeds the capabilities of EMLLMs.
+DDF significantly reduces the amount of training data for Cas-SEAT, lowering its training cost and ensuring its applicability in resource-constrained environments.
+As the first work to explore self-evaluation augmented training, Cas-SEAT and the Cas-SEAT-DDF dataset we constructed provide a valuable foundation and reference for future research in this field.
+Experimental results show that Cas-SEAT improves self-evaluation ability by 19.68%, 55.57%, and 46.79% on the MathVista, Math-V, and We-Math datasets, respectively, significantly outperforming existing methods.
+Furthermore, with the combined effect of Cas-SEAT and DDF, we can use a 7B parameter open-source EMLLMs for data labeling, and the 7B model’s performance exceeds that of a 13B model, fully demonstrating the potential of our method.
 In summary, our contributions are as follows:
-  • We propose SEAT, exploring how EMLLMs improve their CoT reasoning and self-evaluation capabilities by learning from CoT-generated self-evaluation data.
-To the best of our knowledge, this is the first study of its kind on EMLLMs and is rare even among LLMs and MLLMs.
-  • We designed Cas-SEAT to effectively address the challenges EMLLMs face with handling long prompt inputs and lengthy CoT and self-evaluation outputs, successfully enhancing self-evaluation capabilities while maintaining CoT reasoning abilities.
-  • We construct the Cas-SEAT Dataset, the first dataset tailored for self-evaluation augmented training of EMLLMs. It is low-cost yet effective, facilitating future research.
-  • We conducted extensive experiments across multiple datasets, exploring different model architectures (LLaVA-1.5, Qwen2-VL) and parameter scales (7B, 2B).
-We performed a detailed analysis of overall performance, variations in performance across different aspects, and further investigated the CoT reasoning ability and self-evaluation capability.
-The experimental results demonstrate that utilizing Cas-SEAT significantly enhances performance.
-2 Related Work
+  • We analyzed the bottlenecks limiting the self-evaluation ability of EMLLMs and proposed the SEAT method to enhance this ability.
+To our knowledge, this is the first research focused on self-evaluation for EMLLMs.
+  • We designed Cas-SEAT, which effectively addresses the challenges EMLLMs face when handling lengthy prompt inputs and long CoT and self-evaluation outputs.
+It successfully improves self-evaluation ability while maintaining CoT reasoning capability.
+  • We designed DDF and constructed the Cas-SEAT-DDF dataset, the first dataset tailored for self-evaluation augmented training of EMLLMs.
+It is cost-effective and efficient, facilitating future research.
+  • We conducted extensive experiments on multiple datasets, exploring the applicability of our method across different model architectures and parameter sizes.
+The results show that our method not only significantly improves the performance of EMLLMs but also outperforms many models that are much larger than the main model.
+2. Related Work
 CoT Reasoning.
-CoT reasoning Huang et al.
-(2022); Hoffmann et al.
-(2022); Chowdhery et al.
-(2022); Wei et al.
-(2022a); Zhang et al.
-(2022); Kojima et al.
-(2022) enhances the performance of LLMs by forcing them to perform step-by-step reasoning.
-The initial idea was proposed by Wei et al.
-(2022b), demonstrating that simply modifying the prompt could significantly improve model performance on complex tasks.
-Following Wei et al.
-(2022b), numerous studies have shown that LLMs can perform CoT reasoning through zero-shot CoT prompting (Kojima et al.
-(2022)) and few-shot CoT prompting (Wei et al.
-(2022b); Zhang et al.
-(2022)).
-Improvements have been made in areas such as self-consistency (Wang et al., 2022b), least-to-most prompting (Zhou et al., 2022), dynamic least-to-most prompting (Drozdov et al., 2022), bootstrapping training (Zelikman et al., 2022), and self-training (Huang et al., 2022).
-Wang et al.
-(2022b, a) refined the original CoT prompting by marginalizing over diverse reasoning paths, while Zelikman et al.
-(2022); Huang et al.
-(2022) improved CoT prompting through bootstrapping on self-generated CoT prompts.
-Li et al.
-(2022a) introduced a voting classifier to filter sampled CoT prompts before the final prediction.
-Additionally, prompt enhancement and selection (Shum et al., 2023), meta-heuristic approaches (Pan et al., 2023), and meta-graph prompting (Pan et al., 2024) have further advanced standard CoT prompting.
-Existing CoT studies pay little attention to the correctness of synthesized CoT data, especially lacking focus on the CoT capabilities of EMLLMs.
-CoT reasoning is more challenging for EMLLMs.
-Our Cas-SEAT improves the quality of synthesized training data through self-evaluation and enhances the CoT reasoning capabilities of EMLLMs.
+Chain-of-Thought (CoT) reasoning (Huang et al., 2022; Hoffmann et al., 2022; Chowdhery et al., 2022; Wei et al., 2022a; Zhang et al., 2022; Kojima et al., 2022; Wei et al., 2022b) improves the performance of large language models (LLMs) by forcing them to perform step-by-step reasoning.
+Subsequent improvements in CoT reasoning for LLMs have been made in areas such as zero-shot CoT prompting (Kojima et al., 2022), few-shot CoT prompting (Wei et al., 2022b; Zhang et al., 2022), self-consistency (Wang et al., 2022b), multiple reasoning paths Wang et al.
+(2022b, a), minimal-to-maximal prompting (Zhou et al., 2022), dynamic minimal-to-maximal prompting (Drozdov et al., 2022), guided training (Zelikman et al., 2022), and self-training (Zelikman et al., 2022; Huang et al., 2022).
+To optimize CoT prompts, Li et al.
+(2022a) proposed filtering CoT prompts using a voting classifier before the final prediction, (Shum et al., 2023) focused on prompt enhancement and selection, and meta-heuristic methods (Pan et al., 2023) and meta-graph prompts (Pan et al., 2024) have also been employed to further optimize CoT prompting.
+Existing CoT research paid little attention to correctness checking at each step, particularly lacking focus on EMLLMs.
+However, for EMLLMs, CoT reasoning is more prone to shortcuts and mistakes.
+Our Cas-SEAT improves the length and accuracy of CoT reasoning in EMLLMs through self-evaluation.
 LLM Self-Evaluation.
-Recent research on LLM calibration has shown that current LLMs can achieve well-calibrated predictions in specific tasks (Rae et al., 2021; Kadavath et al., 2022; Guo et al., 2017; Kadavath et al., 2022; Jiang et al., 2021; Kuhn et al., 2023; Fu et al., 2023).
-For instance, Liu et al.
-Liu et al.
-(2023a) leveraged CoT prompting to use GPT-4 for text evaluation.
-Kocmi and Federmann Kocmi and Federmann (2023); Xu et al.
-Xu et al.
-(2023) designed methods focused on enabling LLMs to produce more fine-grained correction annotations.
-Additionally, Koo et al.
-Koo et al.
-(2023), Zheng et al.
-Zheng et al.
-(2023), Chang et al.
-Chang et al.
-(2023), Deutsch et al.
-Deutsch et al.
-(2022), and Liu et al.
-Liu et al.
-(2023b) explored issues of self-reinforcing bias and fairness in LLMs.
-It is worth noting that scaling up model size plays a crucial role in improving calibration (Rae et al., 2021; Wei et al., 2022a).
-As LLMs exhibit strong calibration capabilities, more research has focused on prompting LLMs for self-evaluation (Golovneva et al., 2022; Zhang et al., 2023; Shinn et al., 2023; Madaan et al., 2024; Paul et al., 2023).
-Although dedicated evaluators have proven effective in assessing the correctness of reasoning data (Li et al., 2022b; Xu et al., 2024), self-evaluation is more efficient compared to using external models for evaluation, as it eliminates the need for additional annotations and evaluators (Li et al., 2022a).
-Unlike existing methods that optimize the reasoning process through self-evaluation, our approach uses synthetic self-evaluation data to enhance model training.
-Moreover, we focus on EMLLMs, which introduce additional challenges to the training and reasoning processes of self-evaluation.
-3 Methodology
-Figure 2: Overview of the method. It illustrates the prompts designed for Augmented Self-Evaluation and Augmented Cascading Self-Evaluation, along with the corresponding training data generated.
-3.1 Problem Formulation and Notations
-Data and Model.
+Using additional evaluators to assess the correctness of inference data has been proven effective (Li et al., 2022b; Xu et al., 2024; Liu et al., 2023a), but self-evaluation eliminates the need for additional annotations and evaluators (Li et al., 2022a), making it more efficient.
+Currently, LLMs demonstrate strong calibration capabilities, and more and more research focuses on using prompts to enable LLMs to perform self-evaluation (Rae et al., 2021; Kadavath et al., 2022; Guo et al., 2017; Kadavath et al., 2022; Jiang et al., 2021; Kuhn et al., 2023; Fu et al., 2023; Golovneva et al., 2022; Zhang et al., 2023; Shinn et al., 2023; Madaan et al., 2024; Paul et al., 2023).
+Kocmi and Federmann (Kocmi and Federmann, 2023) and Xu et al.
+(Xu et al., 2023) designed methods to guide LLMs to generate more fine-grained corrective annotations.
+In addition, Koo et al.
+(Koo et al., 2023), Zheng et al.
+(Zheng et al., 2023), Chang et al.
+(Chang et al., 2023), Deutsch et al.
+(Deutsch et al., 2022), and Liu et al.
+(Liu et al., 2023b) explored issues of self-amplifying bias and fairness in LLMs.
+Notably, the scaling up of models plays a crucial role in improving calibration capabilities (Rae et al., 2021; Wei et al., 2022a).
+Compared to larger MLLMs, EMLLMs are fragile, with inherently weaker self-evaluation abilities and an inability to handle long token inputs and outputs.
+Moreover, the foundation of self-evaluation—the CoT reasoning ability—is also weak.
+Our Cas-SEAT focuses on these issues and significantly improves the self-evaluation ability of EMLLMs.
+3. Methodology
+Figure 2. Overview of the method. It illustrates the prompts designed for Augmented Self-Evaluation and Augmented Cascading Self-Evaluation, along with the corresponding training data generated.
+3.1. Problem Formulation and Notations
+3.1.1. Data and Model.
 The primary EMLLM model is denoted as $\mathcal{M}$, with parameters $\Theta_{0}$.
-To achieve better performance on the target task, $\mathcal{M}$ needs to be fine-tuned on the dataset $\mathcal{D}$.
-We use $\{\mathcal{X},\mathcal{Y}\}$ to represent a sample, $\mathcal{X}$ include images $\mathcal{I}$ and queries $\mathcal{Q}$, and $\mathcal{Y}$ represents answers.
+To achieve better performance on the target task, $\mathcal{M}$ needs to be fine-tuned on the dataset $\mathcal{D}$, We use $\{\mathcal{X},\mathcal{Y}\}$ to represent a sample, $\mathcal{X}$ include images $\mathcal{I}$ and queries $\mathcal{Q}$, and $\mathcal{Y}$ represents answers.
 After fine-tuning, $\Theta_{0}$ will be optimized to $\Theta$.
@@
 A small subset of $\mathcal{S}_{\rm{cot}}$ is selected, denoted as $\mathcal{S}_{\rm{cot}}^{\rm{sub}}$, and after being evaluated by $\mathcal{G}$, generates cascaded self-evaluation training data, denoted as $\mathcal{S}_{\rm{cse}}$.
-Prompt.
+3.1.2. Prompt.
 The prompts used for $\mathcal{M}$ inference and self-evaluation are denoted as $P$ and $P_{se}$, respectively.
 The prompts used by $\mathcal{G}$ to generate cascading CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{se}$, respectively.
-The prompts used by $\mathcal{G}$ to generate CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{cse}$, respectively.
+The prompts used by $\mathcal{G}$ to generate CoT data and evaluation data are denoted as $P_{\rm{cot}}$ and $P_{\rm{{cse}}}$, respectively.
 Identical symbols indicate identical prompts.
-Formula.
-We formalize the existing studies to facilitate a formula-based comparison of their differences. $l$ represents the loss function.
+3.1.3. Formula.
+We use $l$ to represent the loss function and $\mathcal{L}$ to represent the total loss.
 Inference:
@@
 $\operatorname*{arg\,min}\limits_{\Theta_{0}}\mathcal{L}=\sum\limits_{\mathcal{X},\mathcal{Y}\in\mathcal{D}}l(\mathcal{Y},\mathcal{M}(P\!\!\quad\text{or}\!\!\quad P_{\rm{cot}},\mathcal{X};\Theta_{0}))$
-Augmented CoT:
-$&\mathcal{S}_{\rm{cot}}=\mathcal{G}(P_{\rm{cot}},\mathcal{X};\Theta)\\ &\operatorname*{arg\,min}_{\Theta}\mathcal{L}=\sum_{\mathcal{X},\mathcal{Y}\in\mathcal{S}_{\rm{cot}}}l(\mathcal{Y},\mathcal{M}(P\!\!\quad\text{or}\!\!\quad P_{\rm{cot}},\mathcal{X};\Theta))\end{aligned}\right.$
 SEAT:
@@
 […]
-3.2 SEAT
-Figure 3: Comparison of evaluation in terms of model performance improvement
+3.2. Double-level Data Filtering
+Our data filtering process is conducted in two stages: source data filtering and labeled data filtering.
+The filtering methods include EMLLM filtering and non-EMLLM filtering.
+Samples that are filtered out are collectively referred to as “rejected samples” in this paper.
+In the source data filtering stage, we use the MathV360K dataset (Shi et al., 2024), a commonly used dataset for multimodal large language model training, and apply EMLLM filtering.
+The filtering criteria are as follows: 1) Image quality: Based on image clarity.
+Samples with low image clarity are discarded.
+2) Text quality: Based on the relevance between text and image and whether the question is clear and well-defined.
+Samples with mismatched text and images, or with vague questions, are discarded.
+3) Question domain: Based on whether specialized domain knowledge is required.
+Overly specialized scientific questions, such as medical CT image diagnostics, are removed.
+4) Question difficulty: Based on whether a more powerful model can answer it correctly.
+We randomly sample some samples and use Qwen2-VL-7B to attempt answers.
+If the accuracy of the answers is close to random values, the sample is discarded.
+In the labeled data filtering stage, we use Qwen2-VL-7B to label the data, followed by non-EMLLM filtering based on Qwen’s responses.
+The filtering criteria are as follows: 1) Text quality: Samples with garbled responses or responses in languages other than English are discarded.
+2) Text length: Samples with excessively long responses are discarded.
+3) Text format: The response format is standardized, and samples that do not conform to the format are discarded.
+3.3. Vanilla SEAT
+Figure 3. Comparison of evaluation in terms of model performance improvement
 Since self-evaluation has been proven effective in LLMs and MLLMs, we tested the self-evaluation performance of EMLLM, as shown in Figure 3.
@@
 The learning process can be formulated as Equation 3.
-3.3 Cas-SEAT
+3.4. Cas-SEAT
 Due to the parameter limitations of EMLLM, its performance significantly degrades in scenarios involving long-token inputs and outputs.
@@
 In the above equation, $A:=B$ means assigning the value of $B$ to $A$. Then, $\mathcal{M}$ is fine-tuned on $\mathcal{S}_{\rm{cse}}$ like Equation 4.
-4 Experiments
-We conducted experiments to evaluate the effectiveness and generalizability of Cas-SEAT.
-4.1 Experimental Setup
-4.1.1 Datasets
-We train on MathV360K dataset and evaluate on Math-Vista, Math-V, We-Math.
-MathV360K is a widely used dataset for MLLM training.
-Math-Vista), Math-V, We-Math are three widely used public benchmarks for MLLM evaluation.
-MathVista
-Method Average Multi-choice Free-form Text Integer General VQA Math-targeted VQA
-Base (inference) 0.2850 0.4407 0.1022 0.4407 0.1124 0.3391 0.2389
+Table 1.
+Comparison of Cas-SEAT and the Baselines on MMMU Dataset.
+BUS, TE, AD, HM, SCI and HSS respectively denotes the Business, Tech and Engineering, Art and Design, Health and Medicine, Science, Humanities and Social Science.
+MMMU
+Model Extra Data Total Accuracy BUS TE AD HM SCI HSS
+Heuristics Baselines
+Random Choice Base 0.2210 0.2470 0.2140 0.2920 0.2070 0.1800 0.2000
+Open-Source Models Inference
+miniGPT4-7B Base 0.2680 0.2130 0.2380 0.2920 0.3070 0.2870 0.2920
+CogVLM-17B Base 0.3210 0.2560 0.2890 0.3800 0.3120 0.2510 0.4150
+LLaVA-v1.5-13B Base 0.3640 0.2270 0.3140 0.5170 0.3870 0.2930 0.5330
 […]
-Finetune (inference) 0.3160 0.4481 0.1609 0.4481 0.1770 0.3196 0.3130
-Finetune (evaluation) 0.3490 0.4926 0.1804 0.4926 0.1986 0.3543 0.3444
-CoT (inference) 0.3380 0.4815 0.1696 0.4815 0.1866 0.3326 0.3426
-[…]
-SEAT (inference) 0.2760 0.4278 0.0978 0.4278 0.1077 0.2957 0.2593
-SEAT (evaluation) 0.2850 0.4389 0.1043 0.4389 0.1148 0.3196 0.2556
-Cas-SEAT (inference) 0.3390 0.4889 0.1630 0.4889 0.1794 0.3652 0.3167
-Cas-SEAT (evaluation) 0.4500 0.6222 0.2478 0.6222 0.2727 0.4848 0.4204
-Improve 19.68% 15.07% 31.04% 15.07% 31.04% 11.50% 17.01%
-Table 1: Comparison of Cas-SEAT and the Baseline on MathVista Dataset
-Math-V
-Method All Level1 Level2 Level3 Level4 Level5 ALG ARI CG COM
-Base (inference) 0.0526 0.0800 0.0690 0.0364 0.0444 0.0299 0.0000 0.0000 0.2941 0.0526
+Finetune 0.3559 0.2600 0.3571 0.5133 0.3067 0.2800 0.4333
+CoT 0.3118 0.1800 0.2666 0.4200 0.2800 0.2733 0.4916
+SEAT 0.3204 0.2400 0.2857 0.4066 0.2800 0.2600 0.5000
+LLaVA-v1.5-7B Cas-SEAT 0.3226 0.2667 0.3238 0.3733 0.3400 0.2267 0.4250
+Open-Source Models Self-Evaluation
 […]
@@
 […]
+SEAT 0.3462 0.2600 0.3333 0.4133 0.3066 0.2866 0.5166
+LLaVA-v1.5-7B Cas-SEAT 0.5193 0.4933 0.5000 0.6333 0.4533 0.4266 0.6416
+Improve 31.60% 48.00% 36.35% 11.75% 15.26% 39.09% 22.21%
+Table 2.
+Comparison of Cas-SEAT and the Baselines on MathVista Dataset.
+FQA, MWP, ALG, ARI, LOG, NUM and STA respectively denote figure QA, math word problem, algebraic, arithmetic, logical, numeric, and statistical.
+MathVista
+Model Extra Data Average FQA MWP ALG ARI LOG NUM STA
+Close-Source Models Inference
+Gemini 1.0 Nano 2 Base 0.3060 0.2860 0.3060 0.2710 0.2980 0.1080 0.2080 0.3350
+Gemini 1.0 Pro Base 0.4520 0.4760 0.3920 0.4520 0.3880 0.1080 0.3260 0.5680
+GPT-4V Base 0.4990 0.4310 0.5750 0.5300 0.4900 0.2160 0.2010 0.5580
+Open-Source Models Inference
+InstructBLIP-7B Base 0.2530 0.2310 0.1830 0.2180 0.2710 0.1890 0.2040 0.2310
+LLaVA-13B Base 0.2610 0.2680 0.1610 0.2730 0.2010 0.2430 0.1830 0.2510
+SPHINX-V1-13B Base 0.2750 0.2340 0.2150 0.2560 0.2810 0.1620 0.1740 0.2360
+LLaVA-v1.5-13B Base 0.324 0.2677 0.2366 0.3701 0.272 0.1622 0.2639 0.2558
+[…]
+Finetune 0.3160 0.2416 0.3011 0.3665 0.2890 0.1081 0.2431 0.2724
+CoT 0.3380 0.2416 0.2903 0.4413 0.2805 0.1081 0.2153 0.2658
+SEAT 0.2760 0.1970 0.1667 0.3665 0.2181 0.0811 0.1667 0.2093
+LLaVA-v1.5-7B Cas-SEAT 0.3390 0.3011 0.2581 0.3986 0.2805 0.1622 0.1806 0.3023
+Open-Source Models Self-Evaluation
+Base 0.3530 0.2974 0.1613 0.4021 0.3088 0.1622 0.2500 0.2857
 […]
 […]
-SEAT (inference) 0.0592 0.1132 0.0488 0.0357 0.0000 0.0882 $0$ul0.0526 0.1579 0.0000
-SEAT (evaluation) 0.0888 0.1509 0.0732 0.0714 0.0000 0.1324 $0$ul0.0526 0.1579 0.0526
-Cas-SEAT (inference) 0.1711 0.1321 0.1341 0.2321 0.1778 0.1912 0.1579 0.1579 0.2632 0.1579
-Cas-SEAT (evaluation) 0.2763 0.2642 0.2439 0.2500 0.3111 0.3235 0.3158 0.1579 0.4211 0.2105
-Improve 55.57% 27.33% 25.01% 179.96% 40.01% 57.12% 199.91% 200.19% 14.31% 99.91%
-Table 2: Comparison of Cas-SEAT and the Baseline on Math-V Dataset
+SEAT 0.3490 0.3048 0.2634 0.4413 0.2493 0.2162 0.1736 0.2990
+LLaVA-v1.5-7B Cas-SEAT 0.4500 0.4201 0.4032 0.4733 0.4023 0.4054 0.2986 0.4086
+Improve 19.68% 37.83% 24.98% 6.41% 13.61% 87.51% 7.49% 35.16%
+4. Experiments
+We conducted experiments to evaluate the effectiveness and generalizability of Cas-SEAT.
+4.1. Experimental Setup
+4.1.1. Datasets
+We train on MathV360K (Shi et al., 2024) dataset and evaluate on MMMU (Yue et al., 2024), Math-Vista (Lu et al., 2023), Math-V (Wang et al., 2024b), We-Math (Qiao et al., 2024).
+MathV360K is a widely used dataset for MLLM training.
+MMMU is a widely used public benchmarks for MLLM evaluation and Math-Vista, Math-V, We-Math are three widely used public benchmarks for MLLM evaluation in math domain.
+Table 3. Comparison of Cas-SEAT and the Baselines on We-Math dataset
 We-Math
-Method Avg(Strict) IG(Strict) CM(Strict) Avg(Loose) IG(Loose) CM(Loose)
-Base (inference) 0.0143 0.0171 0.0057 0.0657 0.0171 0.0571
+Model Extra Data Avg(Strict) IG(Strict) CM(Strict) Avg(Loose) IG(Loose) CM(Loose)
+Close-Source Models Inference
+GPT-4V Base 0.3105 0.1448 0.2381 0.5143 0.1448 0.0333
+Gemini 1.5 Pro Base 0.2638 0.1124 0.2076 0.4600 0.1124 0.1203
+Qwen-VL-Max Base 0.1048 0.0762 0.0667 0.2552 0.0762 0.2028
+Open-Source Models Inference
+LLaVA-1.5-13B Base 0.0248 0.0114 0.019 0.0952 0.014 0.0895
+LLaVA-v1.6-13B Base 0.0524 0.0324 0.0362 0.2200 0.0324 0.2621
+DeepSeek-VL-7B Base 0.0629 0.0457 0.0400 0.2095 0.0457 0.2899
+G-LLaVA-13B Base 0.0648 0.0457 0.0419 0.2229 0.0457 0.3598
 […]
-Finetune (inference) 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
-Finetune (evaluation) 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
-CoT (inference) 0.0438 0.0305 0.0286 0.1752 0.0305 0.1600
+Finetune 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
+CoT 0.0438 0.0305 0.0286 0.1752 0.0305 0.1600
+SEAT 0.0105 0.0095 0.0057 0.0429 0.0095 0.0381
+LLaVA-v1.5-7B Cas-SEAT 0.0533 0.0343 0.0362 0.2114 0.0343 0.1943
+Open-Source Models Self-Evaluation
 […]
-SEAT (inference) 0.0105 0.0095 0.0057 0.0429 0.0095 0.0381
-SEAT (evaluation) 0.0114 0.0114 0.0057 0.0495 0.0114 0.0438
-Cas-SEAT (inference) 0.0533 0.0343 0.0362 0.2114 0.0343 0.1943
-Cas-SEAT (evaluation) 0.1076 0.0438 0.0857 0.3495 0.0438 0.3276
+Finetune 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
+[…]
+SEAT 0.0114 0.0114 0.0057 0.0495 0.0114 0.0438
+LLaVA-v1.5-7B Cas-SEAT 0.1076 0.0438 0.0857 0.3495 0.0438 0.3276
 Improve 46.79% 9.50% 55.25% 23.54% 9.50% 24.61%
-Table 3: Comparison of Cas-SEAT and the baselines on We-Math dataset
-4.1.2 Baselines
+Table 4. Comparison of Cas-SEAT and the Baselines across multiple dimensions on the MathVista Dataset.
+MathVista
+Model Extra Data Multi-choice Free-form Text Integer General VQA Math-targeted VQA
+Base 0.4407 0.1022 0.4407 0.1124 0.3391 0.2389
+Finetune 0.4481 0.1609 0.4481 0.1770 0.3196 0.3130
+CoT 0.4815 0.1696 0.4815 0.1866 0.3326 0.3426
+SEAT 0.4278 0.0978 0.4278 0.1077 0.2957 0.2593
+LLaVA-v1.5-7B Cas-SEAT 0.4889 0.1630 0.4889 0.1794 0.3652 0.3167
+Open-Source Models Self-Evaluation
+[…]
+Finetune 0.4926 0.1804 0.4926 0.1986 0.3543 0.3444
+[…]
+SEAT 0.4389 0.1043 0.4389 0.1148 0.3196 0.2556
+LLaVA-v1.5-7B Cas-SEAT 0.6222 0.2478 0.6222 0.2727 0.4848 0.4204
+Improve 15.07% 31.04% 15.07% 31.04% 11.50% 17.01%
+Table 5.
+Detailed comparison of Cas-SEAT and the Baselines on We-Math dataset.
+UCU, AL, CPF, UPF, CSF, USF, and BTF respectively denote Understanding and Conversion of Units, Angles and Length, Calculation of Plane Figures, Understanding of Plane Figures, Calculation of Solid Figures, Understanding of Solid Figures, and Basic Transformations of Figures.
+We-Math
+Model Extra Data UCU AL CPF UPF CSF USF BTF
+Open-Source Models Inference
+[…]
+Finetune 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.1950
+[…]
+SEAT 0.0942 0.0263 0.1099 0.1476 0.1083 0.1624 0.1176
+LLaVA-v1.5-7B Cas-SEAT 0.2679 0.3439 0.3240 0.3216 0.3230 0.2408 0.3240
+Open-Source Models Self-Evaluation
+[…]
+Finetune 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.2172
+[…]
+SEAT 0.0942 0.0263 0.1179 0.1476 0.1130 0.1624 0.1176
+LLaVA-v1.5-7B Cas-SEAT 0.4256 0.1982 0.4294 0.3546 0.3303 0.3431 0.3462
+Improve 8.32% -14.42% 24.36% 16.76% 11.44% 4.60% 34.71%
+Table 6. Comparison on tasks of varying difficulty.
+We-Math Math-V
+Model Extra Data One-step Two-step Three-step All Level1 Level2 Level3 Level4 Level5
+Open-Source Models Inference
+Base 0.1621 0.1472 0.1394 0.0526 0.0800 0.0690 0.0364 0.0444 0.0299
+Finetune 0.3004 0.2750 0.3394 0.1743 0.2075 0.1951 0.0893 0.1778 0.1912
+CoT 0.2905 0.2500 0.2000 0.1414 0.2075 0.1951 0.0893 0.1778 0.1912
+SEAT 0.1210 0.1361 0.1091 0.0757 0.1400 0.0690 0.0364 0.0444 0.0896
+LLaVA-v1.5-7B Cas-SEAT 0.3103 0.2861 0.2364 0.1711 0.2075 0.1951 0.0893 0.1778 0.2059
+Open-Source Models Self-Evaluation
+[…]
+[…]
+[…]
+SEAT 0.1243 0.1472 0.1091 0.1447 0.1509 0.1098 0.0714 0.2222 0.1912
+LLaVA-v1.5-7B Cas-SEAT 0.3909 0.3528 0.4788 0.2763 0.2642 0.2439 0.2500 0.3111 0.3235
+Improve 20.54% 13.40% 27.41% 55.57% 27.33% 25.01% 7.71% 40.01% 57.12%
+4.1.2. Baselines
 To verify the applicability, the following EMLLMs are implemented and compared with the counterparts combined with the proposed method.
-We primarily analyzed the effectiveness of our method based on the EMLLMs LLaVA-v1.5(7B) Liu et al.
-(2024), Qwen2-VL(2B) Wang et al.
-(2024).
+We primarily analyzed the effectiveness of our method based on the EMLLMs LLaVA-v1.5(7B) (Liu et al., 2024a), Qwen2-VL(2B) (Wang et al., 2024a).
 Since current EMLLMs research often involves training on various datasets, including those frequently used as test datasets, such as the ones we selected, and many MLLM studies do not disclose their training datasets, we opted not to include EMLLMs published after the release of these datasets to ensure a fair comparison.
-4.1.3 Implementation Details
-“Base”: The pretrained MLLM Liu et al.
-(2024); Wang et al.
-(2024).
-“Finetune”: The pretrained MLLM is lora fine-tuned on Math360k Hu et al.
-(2022).
+We also included the following models as references: miniGPT4 (Zhu et al., 2023), CogVLM (Wang et al., 2023a), LLaVA-v1.6 (Liu et al., 2024b), Gemini 1.0 (Team et al., 2023), Gemini 1.5 (Team et al., 2023), GPT-4V (OpenAI, 2023), Shared GPT-4V (Chen et al., 2024), SPHINX-V1 (Lin et al., 2023), G-LLaVA (Gao et al., 2023), DeepSeek-VL (Lu et al., 2024), Qwen-VL (Bai et al., 2023).
+4.1.3. Implementation Details
+“Base”: The pretrained MLLM (Liu et al., 2024a; Wang et al., 2024a).
+“Finetune”: The pretrained MLLM is lora fine-tuned on Math360k (Hu et al., 2022).
 “CoT”: Qwen2-VL (7B) performs CoT reasoning on Math360k.
-The MLLM then learns from this data Wang et al.
-(2023); Wadhwa et al.
-(2024).
+The MLLM then learns from this data (Wang et al., 2023b; Wadhwa et al., 2024).
 “SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k.
-Qwen2-VL (7B) evaluates the reasoning outputs of the MLLM, and the MLLM learns from this evaluated data Xiong et al.
-(2024).
+Qwen2-VL (7B) evaluates the reasoning outputs of the MLLM, and the MLLM learns from this evaluated data (Xiong et al., 2024).
 “Cas-SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k.
@@
 These evaluated data are combined with CoT data, and the MLLM learns from this combined dataset.
-We-Math
-Method UCU AL CPF UPF CSF USF BTF One-step Two-step Three-step
-[…]
-[…]
-Finetune (inference) 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.1950 0.3004 0.2750 0.3394
-Finetune (evaluation) 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.2172 0.3021 0.2750 0.3394
-[…]
-[…]
-SEAT (inference) 0.0942 0.0263 0.1099 0.1476 0.1083 0.1624 0.1176 0.1210 0.1361 0.1091
-SEAT (evaluation) 0.0942 0.0263 0.1179 0.1476 0.1130 0.1624 0.1176 0.1243 0.1472 0.1091
-Cas-SEAT (inference) 0.2679 0.3439 0.3240 0.3216 0.3230 0.2408 0.3240 0.3103 0.2861 0.2364
-Cas-SEAT (evaluation) 0.4256 0.1982 0.4294 0.3546 0.3303 0.3431 0.3462 0.3909 0.3528 0.4788
-Improve 8.32% -14.42% 24.36% 16.76% 11.44% 4.60% 34.71% 20.54% 13.40% 27.41%
-Table 4: Detailed comparison of Cas-SEAT and the baselines on We-Math dataset
 After obtaining outputs on the test dataset, we first use a more powerful MLLM to extract answers for the MathVista dataset, followed by using regular expressions to calculate accuracy.
 For Math-V and We-Math, we directly use the regular expressions provided in the datasets to calculate accuracy.
-4.2 Experimental Results
-4.2.1 Overall Assessment of Cas-SEAT
-As shown in Tables 1, 2, 3, and 4, we analyzed the experimental results on the MathVista, Math-V, and We-Math datasets.
+Figure 4.
+A bar chart analysis sample in MathVista.
+Green background indicates the raw data, red text represents incorrect reasoning processes (sometimes with no reasoning process), pink background and yellow background denote results from direct reasoning and self-evaluation, respectively.
+Blue text and blue background indicate the corrected reasoning process and corrected results, respectively.
+4.2. Experimental Results
+4.2.1. Overall Assessment of Cas-SEAT
+As shown in Tables 1, 2, 3, 4, 5, 6, we analyze the experimental results on the MMMU, MathVista, Math-V, and We-Math datasets.
 We evaluated the accuracy of two types of outputs for each method: direct inference (inference) and self-evaluation (evaluation).
-In addition to overall performance, we conducted tests on multiple subsets of the three datasets, including generalization capability, data types, problem difficulty, and task types.
+In addition to overall performance, we conducted tests on multiple subsets of these datasets, including generalization capability, data types, problem difficulty, and task types.
 Below is the correspondence between these categories and the columns in the tables.
-(Due to dataset limitations, some categories could not be tested for certain datasets.): Overall Performance: MathVista (Table 1: Average), Math-V (Table 2: All), We-Math (Table 3: Score (Strict), Score (Loose)).
-Cas-SEAT shows very significant improvements over the baseline on all datasets, achieving around a 50% improvement on both Math-V and We-Math.
-Generalization Capability: We-Math (Table 3: Inadequate Generalization (Strict), Complete Mastery (Strict), Inadequate Generalization (Loose), Complete Mastery (Loose)).
-The improvement range of 9.5% to 55.25% demonstrates Cas-SEAT’s strong generalization capabilities, excelling at problems involving more integrated knowledge points, as evidenced by the significant improvement in Complete Mastery (CM).
+(Due to dataset limitations, some categories could not be tested for certain datasets.): Overall Performance: MMMU (Table 1: Total Accuracy), MathVista (Table 2: Average), We-Math (Table 3: Score (Strict), Score (Loose)).
+Cas-SEAT shows very significant improvements over the baseline, achieving around a 20% $\sim$ 50% improvement on both the aforementioned datasets, which validates the effectiveness of the proposed approach.
+Furthermore, we conducted a detailed comparison of the performance under different Data Types, Problem Difficulty, and Task Types.
 Data Types: MathVista (Table 4: Text, Integer).
 Cas-SEAT is better at answering integer-type questions.
-Problem Difficulty: MathVista (Table 1: Multi-choice, Free-form), Math-V (Table 2: Level1, Level2, Level3, Level4, Level5), We-Math (Table 3: One-step, Two-step, Three-step).
-The Math-V subsets are divided into five difficulty levels, with higher levels indicating more complex problems.
+Problem Difficulty: As shown in Table 6, the Math-V subsets are divided into five difficulty levels, with higher levels indicating more complex problems.
 The We-Math subsets are classified according to the number of reasoning steps, with more steps indicating higher difficulty.
 Cas-SEAT shows a more pronounced advantage on more challenging problems.
-Task Types: MathVista: General VQA, Math-targeted VQA; Math-V: Algebra (ALG), Arithmetic (ARI), Combinatorial Geometry (CG), Combinatorics (COM); We-Math: Understanding and Conversion of Units (UCU), Angles and Length (AL), Calculation of Plane Figures (CPF), Understanding of Plane Figures (UPF), Calculation of Solid Figures (CSF), Understanding of Solid Figures (USF), Basic Transformations of Figures (BTF).
-Cas-SEAT exhibits significant improvements over the best baseline method on almost all problem types, especially for numerical computation tasks (Math-V: ALG, ARI).
+Task Types: MMMU (Table 1): BUS, TE, AD, HM, SCI, HSS; MathVista (Table LABEL:tab:main_mathvista,tab:main_mathvista_2): General VQA, Math-targeted VQA, FQA, MWP, ALG, ARI, LOG, NUM, STA; We-Math (Table 5): UCU, AL, CPF, UPF, CSF, USF, BTF.
+Math-V (Table 11 in Appendix): ALG, ARI, CG, COM; Cas-SEAT exhibits significant improvements over the best baseline method on almost all problem types, especially for numerical computation tasks (Math-V: ALG, ARI).
 For VQA tasks, Cas-SEAT performs better on the more difficult Math-targeted VQA.
-In summary, we found that: 1.
+In summary, we found that: (1).
 After fine-tuning, the inference ability of EMLLMs improves, but their self-evaluation ability declines.
@@
 Using synthetic CoT data based on mathematical datasets for enhanced training shows a similar pattern.
-2.
+(2).
 SEAT, which integrates CoT inference and self-evaluation training, suffers a significant drop in CoT inference ability, without a substantial improvement in self-evaluation.
@@
 This demonstrates the detrimental effect of long prompts and outputs on EMLLMs, and the difficulty of maintaining both CoT inference and self-evaluation capabilities simultaneously.
-3.
+(3).
 Cas-SEAT, by separating the CoT inference and self-evaluation tasks and shortening the prompt, preserves robust inference ability.
@@
 The detailed analysis shows that Cas-SEAT is especially adept at more difficult numerical computation tasks.
-4.2.2 Evaluation on Different EMLLMs
-Figure 4: Comparison of Cas-SEAT and the Baseline based on Qwen2-VL(2B).
-Figure 5: A bar chart analysis sample in MathVista.
-Green background indicates the raw data, red text represents incorrect reasoning processes (sometimes with no reasoning process), pink background and yellow background denote results from direct reasoning and self-evaluation, respectively.
-Blue text and blue background indicate the corrected reasoning process and corrected results, respectively.
-As shown Figure 4, we used Qwen2-VL (2B) on Math-V to compare the performance of Cas-SEAT with that of the baseline.
-We take the maximum value of all methods as the circumference and 0 as the center of the circle.
-The results demonstrated a significant advantage for Cas-SEAT over the baselines.
-This indicates that our approach is also applicable to smaller EMLLMs with different architectures.
-4.2.3 Case Study
+4.2.2. Case Study
 We present two representative examples in Figure 4 and Figure 6 (in appendix).
@@
 These conclusions hold consistently across several datasets.
-5 Conclusion
+5. Conclusion
 We proposed SEAT and its enhanced variant, Cas-SEAT, to improve the self-evaluation capabilities of Efficient Multimodal Large Language Models (EMLLMs).
@@
 Experiments show significant gains on benchmark datasets, and the Cas-SEAT Dataset provides a valuable resource for future research, laying a strong foundation for advancing self-evaluation in EMLLMs.
-Limitation
-Potential limitations may include a small amount of resource consumption. The method is designed specifically for EMLLMs, and it may not currently be the optimal solution for larger MLLMs.
 Appendix A Appendix
-A.1 Supplementary Experiments
-A.1.1 Case Study
-Figure 6: A table analysis sample in MathVista. The meanings of the colors are the same as those in Figure 5.
-Figure 6 is a supplementary case, the observed findings can be referred to the section 4.2.3.
-A.1.2 Datasets
+A.1. Supplementary Experiments
+A.1.1. Cost Analysis
+As shown in Table 9, DDF significantly reduces the training data size and consequently the training duration while using the same LLaVA-v1.5-7B model.
+Our method makes LLaVA-v1.5-7B far outperform LLaVA-v1.5-13B.
+In this context, we further compared the inference cost of LLaVA-v1.5-7B under our method with that of LLaVA-v1.5-13B, demonstrating the efficiency of our approach.
+A.1.2. Datasets
 The statistics of the training dataset used in the experiments is shown in Table 10.
-Table 5: Statistics of Datasets.
-Dataset MathV360k CoT Cas-SEAT
-#Samples 339k 160k 167k
-A.1.3 Hyperparameters and Training Schedules
+A.1.3. Hyperparameters and Training Schedules
 We summarize the hyperparameters and training schedules of the LLaVA-1.5 and Qwen2-VL used in the experiments.
@@
 Table 8 shows the settings of the Qwen2-VL training.
-Table 6: Hyperparameters and training schedules of LLaVA.
+Table 7. Hyperparameters of LLaVA-1.5.
 MLLM Hyperparameter Setting
-LLaVA-1.5- 7-hf GPU Tesla A100 (40GB)
+LLaVA-1.5-7B GPU Tesla A100 (40GB)
 Batch size 20
@@
 Warmup steps 50
-Learning tate 2e-5
+Learning rate 2e-5
 Epochs 2
-Table 7: Hyperparameters and training schedules of Qwen2-VL.
+Table 8. Hyperparameters of Qwen2-VL.
 MLLM Hyperparameter Setting
-Qwen2-VL- 2B-Instruct GPU Tesla A100 (40GB)
+Qwen2-VL-2B GPU Tesla A100 (40GB)
 Batch size 20
@@
 Warmup steps 50
-Learning tate 2e-5
+Learning rate 2e-5
 Epochs 2
+Table 9. Comparison of training and inference cost.
+Training speed
+On raw dataset 18h
+On Cas-SEAT-DDF dataset 10h
+Inference speed
+LLaVA-v1.5-7B Input: 7700 tokens/s Output: 100 tokens/s
+LLaVA-v1.5-13B Input: 3900 tokens/s Output: 50 tokens/s
+Table 10. Statistics of Datasets.
+Dataset MathV360k CoT Cas-SEAT
+#Samples 339k 160k 167k
+Table 11. Comparison of Cas-SEAT and the Baselines on Math-V Dataset. ALG, ARI, CG, and COM respectively denote Algebra, Arithmetic, Combinatorial Geometry, and Combinatorics.
+Math-V
+Model Extra Data All Level1 Level2 Level3 Level4 Level5 ALG ARI CG COM
+Heuristics Baselines
+Random Choice Base 0.0717 - - - - - 0.0150 0.0710 0.0970 0.0480
+Close-source Models Inference
+Qwen-VL-Plus Base 0.1072 - - - - - 0.1130 0.1430 0.1270 0.0480
+Qwen-VL-Max Base 0.1559 - - - - - 0.1070 0.2000 0.1690 0.1250
+GeminiPro Base 0.1766 - - - - - 0.1510 0.2070 0.2010 0.1190
+GPT4V Base 0.2276 - - - - - 0.2730 0.3570 0.2110 0.1670
+Open-Source Models Inference
+SPHINX(V2) Base 0.0970 - - - - - 0.0670 0.1290 0.0750 0.0770
+ShareGPT4V-7B Base 0.1053 - - - - - 0.0550 0.1290 0.1010 0.0480
+LLaVA-v1.5-13B Base 0.1112 - - - - -
+[…]
+Finetune 0.1743 0.2075 0.1951 0.0893 0.1778 0.1912 0.1053 0.0000 0.2632 0.0000
+CoT 0.1414 0.2075 0.1951 0.0893 0.1778 0.1912 0.0526 0.0000 0.3684 0.0526
+SEAT 0.0757 0.1400 0.0690 0.0364 0.0444 0.0896 0.1053 0.0000 0.2632 0.0000
+[…]
+Open-Source Models Self-Evaluation
+Base 0.0757 0.1509 0.1098 0.0714 0.2222 0.1765 0.0000 0.0000 0.3529 0.1053
+Finetune 0.1776 0.2075 0.1951 0.0893 0.1778 0.2059 0.1053 0.0000 0.2632 0.0000
+CoT 0.1447 0.1321 0.1341 0.2321 0.1778 0.1912 0.1053 0.0000 0.3684 0.0526
+[…]
+LLaVA-v1.5-7B Cas-SEAT 0.2763 0.2642 0.2439 0.2500 0.3111 0.3235 0.3158 0.1579 0.4211 0.2105
+Improve 55.57% 27.33% 25.01% 7.71% 40.01% 57.12% 100.00% 0.00% 14.31% 33.31%
+A.1.4. Evaluation on Different EMLLMs
+Figure 5. Comparison of Cas-SEAT and the Baseline based on Qwen2-VL(2B).
+As shown Figure 5, we used Qwen2-VL (2B) on Math-V to compare the performance of Cas-SEAT with that of the baseline.
+We take the maximum value of all methods as the circumference and 0 as the center of the circle.
+The results demonstrated a significant advantage for Cas-SEAT over the baselines.
+This indicates that our approach is also applicable to smaller EMLLMs with different architectures.
+A.1.5. Case Study
+Figure 6 is a supplementary case, the observed findings can be referred to the section 4.2.2. Figure 7 is a reject sample case, which is detrimental to the training of EMLLMs.
+Figure 6. A table analysis sample in MathVista. The meanings of the colors are the same as those in Figure 4.
+Figure 7. A case of the reject samples.